# Scalable tensor network algorithm for quantum impurity problems

### v1

@@ I Introduction
The continuous-time quantum Monte Carlo (CTQMC) methods have been the golden standard for quantum impurity solvers [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. They can yield numerically exact results, and are highly efficient on the imaginary-time axis, especially when the sign problem is not serious. However, real-time calculations remain a great challenge for CTMQC. A number of alternatives are also available, which complement CTQMC on the real-time axis or at zero temperature, including exact diagonalization [15, 16, 17, 18, 19, 20, 21, 22], numerical renormalization group [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], hierarchical equation of motion [36, 37, 38, 39, 40, 41, 42, 43], time-evolving matrix product state (MPS) [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Unlike CTQMC, these methods generally suffer from various sources of uncontrolled numerical errors, which could prohibit them from being used as general-purpose quantum impurity solvers.
The Grassmann time-evolving matrix product operator (GTEMPO) method, proposed by some of us, is a promising candidate as a general-purpose quantum impurity solver [55]. GTEMPO represents the integrand of the impurity path integral (PI), referred to as the augmented density tensor (ADT), as a Grassmann MPS (GMPS), in analogous to the time-evolving matrix product operator (TEMPO) method for bosonic impurity problems [56]. As the bath degrees of freedom are integrated out before hand via the Feynman-Vernon influence functional (IF) [57], GTEMPO is naturally free of the bath discretization error, similar to CTQMC. In fact, there are essentially only two sources of numerical errors in GTEMPO, one from discretization of the PI, the other from the MPS bond truncation. The first could be suppressed by using a finer time step size, while the second can be suppressed by using a larger bond dimension for MPS in principle. Meanwhile, as the core algorithms of GTEMPO only rely on the analytical expression of the IF, and that MPS algorithms are naturally free of the sign problem, GTEMPO can be flexibly applied on the imaginary-time axis [58], real-time axis (i.e., Keldysh contour) [59, 60], or even the L-shaped Kadanoff contour [61, 62].
@@
In this work we propose a multi-flavor extension of the GTEMPO method to resolve this issue. As an impurity solver in DMFT, the central observables to calculate are the single-particle Green’s functions which depends at most on two flavors. Therefore the degrees of freedom of the rest flavors can be integrated out before hand, similar to the treatment of the bath (although the flavor degrees of freedom can only be traced out numerically in general). Even though the full ADT may not be compressed, it is highly possible that the reduced ADT, which keeps only the information of the involved flavors, can be largely compressed without significant loss of accuracy. This idea is implemented in this work on top of GTEMPO, referred to as the multi-flavor GTEMPO method in the following. We note that a similar idea has been implemented in the bosonic case, based on TEMPO [64]. We apply this method to study imaginary-time evolution of QIPs with up to three orbitals ($6$ flavors), and benchmark our results against CTQMC. Our numerical results show that this method is indeed effective: under a commonly used semi-circular spectrum function of the bath, we observe that we can obtain highly accurate results (the error is of the same order with the chosen time step size) with bond dimensions $200$ and $700$ in the two-orbital and three-orbital cases respectively, which are far smaller than those from exponential growth. These results illustrate the power of the multi-flavor GTEMPO method to solve large-scale QIPs.
@@ II.2 The multi-flavor GTEMPO method
where $Z_{\mathrm{bath}}$ is the partition function of the bath, $\bar{\bm{a}}_{p}=\{\bar{a}_{p}(\tau)\}$, $\bm{a}_{p}=\{a_{p}(\tau)\}$ are the Grassmann trajectories for the $p$th flavor over the whole imaginary-time interval $[0,\beta]$ ($\beta$ is the inverse temperature), and $\bar{\bm{a}}=\{\bar{\bm{a}}_{p},\bar{\bm{a}}_{q},\cdots\}$, $\bm{a}=\{\bm{a}_{p},\bm{a}_{q},\cdots\}$. The measure $\mathcal{D}[\bar{\bm{a}}\bm{a}]$ is
@@
The function $\Delta(\tau^{\prime},\tau^{\prime\prime})$ in Eq.(8) is usually referred to as the hybridization function, which directly determines the IF, and can be calculated with the spectrum function and the free bath Matsubara Green’s function [58]. The integrand in Eq.(5) defines the augmented density tensor, denoted as
@@
which contains all the information of the multi-time impurity dynamics.
Figure 1: (a) Schematical illustration of the zipup algorithm to to calculate the partition function for an impurity problem with $2$ flavors, where the quasi-2D tensor network is contracted (integrating out the pairs of conjugate Grassmann variables) from left to right and the augmented density tensor $\mathcal{A}$ is calculated on the fly as indicated by the dashed $\times$. (b) The scheme used in the multi-flavor GTEMPO method to calculate the partition function based on the reduced ADT $\mathcal{A}_{2}$ for the second flavor, in which one first calculate $\mathcal{K}\mathcal{I}_{\overline{1}}$ by multiplying $\mathcal{K}$ and $\mathcal{I}_{1}$ and integrating out the first flavor (b1), and then multiply $\mathcal{K}\mathcal{I}_{\overline{1}}$ and $\mathcal{I}_{2}$ to obtain $\mathcal{A}_{2}$ (b2). Again the second multiplication is only performed on the fly similar to (a). The empty circles in (b1) mean that these Grassmann variables do not exist in the corresponding GMPS, while the gray box means that the Grassmann variables inside it will be integrated out after multiplication.
@@ III.1 Toulouse model
We first consider the single-flavor case with $\hat{H}_{\rm imp}=\epsilon_{d}\hat{a}^{\dagger}\hat{a}$, referred to as the Toulouse model, which can be analytically solved [71]. We will fix $\epsilon_{d}=1$. For this model there is only a single $\mathcal{I}_{1}$ and there is no need to use the multi-flavor GTEMPO method. The purpose of this example is merely to find the proper values of the two hyperparameters $\chi$ and $\delta\tau$ that will be used in later multi-flavors calculations.
@@
In Fig. 3(a,b), we study the accuracy of GTEMPO results against different $\chi$ for the Toulouse model at $\beta=10$ and $\beta=50$, respectively. We can see that the GTEMPO results generally agree very well with the analytical solutions. From the left insets, we can see that the mean error goes down monotonically with $\chi$ in both cases. From the insets of both panels, we can see that the errors are generally larger at $\beta=50$ compared to $\beta=10$, even though larger $\chi$s have been used, which indicates that GTEMPO has a larger computational cost at lower temperature.
@@ III.3 Multi-orbital quantum impurity models
for these two cases respectively [72, 73]. In our numerical simulations, we fix $U=4$ and $J=1$, and $\epsilon_{d}$ is set by the half-filling conditions.
@@ III.4 DMFT iterations on the Bethe lattice
Finally, we apply our multi-flavor GTEMPO as an impurity solver in DMFT, to solve the multi-orbital Slater-Kanamori mode [11] on the Bethe lattice. The hopping parameter of the Bethe lattice is chosen to be $1$. To benchmark our results against CTQMC in a step-by-step fashion, we simply choose Eq.(4) as the initial spectral function for both solvers. We have performed $10$ DMFT iterations in the two-orbital case with $\chi_{2}=200$, and $5$ DMFT iterations in the three-orbital case with $\chi_{2}=700$. In both case the DMFT iterations have well converged, with mean errors between the last two iterations to be $5.5\times 10^{-5}$ and $3.7\times 10^{-4}$. The results of the two cases are plotted in Fig. 6(a,b) respectively. In the insets we show the absolute errors between multi-flavor GTEMPO and CTQMC results. We can see that the absolute errors in both cases, and at different iterations, are all of the order $10^{-2}$ or less, and interestingly, the errors in the last DMFT iteration is the smallest in both cases. These results thus illustrate the power of the multi-flavor GTEMPO method as a scalable quantum impurity solver.
@@ Acknowledgements.
Z. L. is partially supported by NSFC (22393913) and the robotic AI-Scientist platform of Chinese Academy of Sciences R. C. acknowledges support from National Natural Science Foundation of China under Grant No. 12104328.

### v2

@@ I Introduction
The continuous-time quantum Monte Carlo (CTQMC) methods have been the golden standard for quantum impurity solvers [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. They can yield numerically exact results, and are highly efficient on the imaginary-time axis, especially when the sign problem is not serious. However, real-time calculations remain a great challenge for CTMQC. A number of alternatives are also available, which complement CTQMC on the real-time axis or at zero temperature, including exact diagonalization [15, 16, 17, 18, 19, 20, 21, 22], numerical renormalization group [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], hierarchical equation of motion (HEOM) [36, 37, 38, 39, 40, 41, 42, 43], time-evolving matrix product state (MPS) [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Unlike CTQMC, these methods generally suffer from various sources of uncontrolled numerical errors, which could prohibit them from being used as general-purpose quantum impurity solvers. For example, the time-evolving MPS method unavoidably suffers from the bath discretization error as the bath is explicitly taken into account, it is also difficult to treat finite temperature baths [53], or scale up to large impurities due to the fast entanglement growth in the impurity-bath wave function [47]. HEOM, in comparison, is free of the bath discretization error, but it truncates the coupled operator equations to a certain order which could result in uncontrolled error, and its computational cost grows exponentially with the order. MPS has also been used in a development of HEOM to avoid the exponential cost growth [41], but its potential to solve large impurity problems is yet to be demonstrated.
[…]
@@
In this work we propose a multi-flavor extension of the GTEMPO method to resolve this issue. As an impurity solver in DMFT, the central observables to calculate are the single-particle Green’s functions which depends at most on two flavors. Therefore the degrees of freedom of the rest flavors can be integrated out before hand, similar to the treatment of the bath (although the flavor degrees of freedom can only be traced out numerically in general). Even though the full ADT may not be compressed, it is highly possible that the reduced ADT, which keeps only the information of the involved flavors, can be largely compressed without significant loss of accuracy. This idea is implemented in this work on top of GTEMPO, referred to as the multi-flavor GTEMPO method in the following. We note that a similar idea has been implemented in the bosonic case, based on TEMPO [66]. We apply this method to study imaginary-time evolution of QIPs with up to three orbitals (i.e., $6$ flavors, as a single electron orbital contains two flavors: spin up and down), and benchmark our results against CTQMC. Our numerical results show that this method is indeed effective: under a commonly used semi-circular spectrum function of the bath, we observe that we can obtain highly accurate results (the error is of the same order with the chosen time step size) with bond dimensions $200$ and $700$ in the two-orbital and three-orbital cases respectively, which are far smaller than those from exponential growth. These results illustrate the power of the multi-flavor GTEMPO method to solve large-scale QIPs.
To this end, we stress that although GTEMPO and the conventional time-evolving MPS methods both make use of MPS and are free of the sign problem, they are sharply different: the former represents the multi-time degrees of freedom of the impurity on the time axis, i.e., the ADT, as an MPS, while the latter represents the impurity-bath wave function at a certain time as an MPS. Therefore GTEMPO does not suffer from bath discretization error and is easy to treat finite temperature baths, GTEMPO is also likely to be more scalable to large impurity problems, especially with the development of the current work.
@@ II.2 The multi-flavor GTEMPO method
where $Z_{\mathrm{bath}}$ is the partition function of the bath, $\bar{\bm{a}}_{p}=\{\bar{a}_{p}(\tau)\}$, $\bm{a}_{p}=\{a_{p}(\tau)\}$ are the Grassmann trajectories for the $p$th flavor over the whole imaginary-time interval $0,\beta]$ ($\beta$ is the inverse temperature), and $\bar{\bm{a}}=\{\bar{\bm{a}}_{p},\bar{\bm{a}}_{q},\cdots\}$, $\bm{a}=\{\bm{a}_{p},\bm{a}_{q},\cdots\}$. The measure $\mathcal{D}[\bar{\bm{a}}\bm{a}]$ is
@@
The function $\Delta(\tau^{\prime},\tau^{\prime\prime})$ in Eq.(8) is usually referred to as the hybridization function, which directly determines the IF, and can be calculated as [58]:
[…]
where $D_{\epsilon}(\tau^{\prime},\tau^{\prime\prime})$ is the free bath Matsubara Green’s function defined as [68]
$D_{\epsilon}(\tau^{\prime},\tau^{\prime\prime})=-[\Theta(\tau^{\prime}-\tau^{\prime\prime})-n(\epsilon)]e^{-\epsilon(\tau^{\prime}-\tau^{\prime\prime})},$ (10)
with $\Theta$ the Heaviside step function and $n(\epsilon)=(e^{\beta\epsilon}+1)^{-1}$ the Fermi–Dirac distribution. The integrand in Eq.(5) defines the augmented density tensor, denoted as
@@
which contains all the information of the multi-time impurity dynamics. Here we point out that the fermionic ADT is essentially the same as a recently proposed concept: the process tensor [69, 70, 71, 72], except that the ADT is often discussed in the context of impurity problems.
Figure 1: (a) Schematical illustration of the zipup algorithm to to calculate the partition function for an impurity problem with 22 flavors, where the quasi-2D tensor network is contracted (integrating out the pairs of conjugate Grassmann variables) from left to right and the augmented density tensor $\mathcal{A}$ is calculated on the fly as indicated by the dashed $\times$. (b) The scheme used in the multi-flavor GTEMPO method to calculate the partition function based on the reduced ADT $\mathcal{A}_{2}$ for the second flavor, in which one first calculate $\mathcal{K}\mathcal{I}_{\overline{1}}$ by multiplying $\mathcal{K}$ and $\mathcal{I}_{1}$ and integrating out the first flavor (b1), and then multiply $\mathcal{K}\mathcal{I}_{\overline{1}}$ and $\mathcal{I}_{2}$ to obtain $\mathcal{A}_{2}$ (b2). Again the second multiplication is only performed on the fly similar to (a). The empty circles in (b1) mean that these Grassmann variables do not exist in the corresponding GMPS, while the gray box means that the Grassmann variables inside it will be integrated out after multiplication.
@@ III.1 Toulouse model
We first consider the single-flavor case with $\hat{H}_{\rm imp}=\epsilon_{d}\hat{a}^{\dagger}\hat{a}$, referred to as the Toulouse model, which can be analytically solved [78]. We will fix $\epsilon_{d}=1$. For this model there is only a single $\mathcal{I}_{1}$ and there is no need to use the multi-flavor GTEMPO method. The purpose of this section is merely to find the proper values of the two hyperparameters $\chi$ and $\delta\tau$, which will be used in the multi-flavor GTEMPO calculations in later sections.
@@
[…]
@@ III.3 Multi-orbital quantum impurity models
for these two cases respectively [79, 80]. In our numerical simulations, we fix $U=4$ and $J=1$, and $\epsilon_{d}$ is set by the half-filling conditions. Here we note that with the spectrum function in Eq.(4), we have chosen a bandwidth $4$ and a coupling strength between the impurity and bath to be $1$. As the energy scale of usual materials is eV, if we take the coupling strength as $1$ eV, which is about $11605$ K, then $\beta=10$ corresponds to $\sim 1160$ K, which is about $4$ times higher than room temperature.
@@
In fact, in Ref. [58], similar two- and three-orbital impurity models have been studied using vanilla GTEMPO, but only $\beta=10$ and $\beta=2$ have been reached for these two cases respectively due to the exponential growth of cost. With multi-flavor GTEMPO, we still expect an exponential growth of $\chi_{2}$ with the number of flavors for general impurity problems, but of the form $O(2^{n}\chi)$ instead of $O(2^{n}\chi^{n})$. The reason is that in multi-flavor GTEMPO, we believe the influence of the rest flavors on the current flavor, after been integrated out, is similar to adding in some effective interaction among the flavors, and if this is the case, the largest contribution it can have on the bond dimension of the ADT is $O(2^{n})$ (we note that in the bosonic version of this method, a large impurity with $21$ spins has been studied [66]). Under this scaling, we would expect $\chi_{2}\approx 3000$ and $\chi_{2}\approx 10000$ are sufficient for the four- and five-orbital cases respectively, under the same $\beta$. In addition, it has been observed in Ref. [58] that the bond dimension $\chi$ of $\mathcal{I}$ grows approximately linearly with the inverse temperature. Therefore we would also expect a linear growth of $\chi_{2}$ for multi-flavor GTEMPO against $\beta$.
@@ III.4 DMFT iterations on the Bethe lattice
Finally, we apply our multi-flavor GTEMPO as an impurity solver in DMFT, to solve the multi-orbital Slater-Kanamori model [11] on the Bethe lattice. The hopping parameter of the Bethe lattice is chosen to be $1$. To benchmark our results against CTQMC in a step-by-step fashion, we simply choose Eq.(4) as the initial spectral function for both solvers. We have performed $10$ DMFT iterations in the two-orbital case with $\chi_{2}=200$, and $5$ DMFT iterations in the three-orbital case with $\chi_{2}=700$. In both case the DMFT iterations have well converged, with mean errors between the last two iterations to be $5.5\times 10^{-5}$ and $3.7\times 10^{-4}$. The results of the two cases are plotted in Fig. 6(a,b) respectively. In the insets we show the absolute errors between multi-flavor GTEMPO and CTQMC results. We can see that the absolute errors in both cases, and at different iterations, are all of the order $10^{-2}$ or less, and interestingly, the errors in the last DMFT iteration is the smallest in both cases. These results thus illustrate the power of the multi-flavor GTEMPO method as a scalable quantum impurity solver.
To this end, we briefly discuss the errors in multi-flavor GTEMPO as well as in general DMFT applications. The first source of error in multi-flavor GTEMPO is the time discretization error, determined by the discrete time step size, e.g., $\delta\tau$ for imaginary-time calculations. In our numerical simulations we have mostly used $\delta\tau=0.05$ and observed an absolute error of the order $10^{-2}$. The time discretization error can be systematically suppressed by using a smaller time step size, or using higher-order Trotter decompositions of the PI [81, 82] or even continuous MPS [83]. The second source of error in multi-flavor GTEMPO is the MPS bond truncation error, determined by $\chi$ and $\chi_{2}$. The MPS bond truncation error can be systematically suppressed by using larger $\chi$ and $\chi_{2}$. As a common practice in MPS algorithms [84], one could even perform an extrapolation against $\chi$ and $\chi_{2}$ to the infinite bond dimension limit to eliminate this error. On the other hand, the DMFT procedure is often tolerant of small errors. For example, in early times the Hirsch-Fye algorithm was widely adopted as an impurity solver for DMFT applications [85], which, besides the Monte Carlo sampling error, has a time discretization error of the order $\delta\tau^{2}$. The time step size $\delta\tau$ in this method is usually of the order $0.1$, and thus the discretization error is of order $10^{-2}$, similar to that in our simulations. In addition, CTQMC is widely adopted as a modern impurity solver [11]. When applied to nonequilibrium DMFT, its error can also reach the order of $10^{-2}$, for example, see Fig. 25 in Ref. [86]. Overall, the errors in GTEMPO can be well controlled and systematically suppressed, compared to its alternatives, but as an impurity solver in DMFT, one may need to balance the numerical accuracy and computational costs, especially for large-scale impurity problems.
@@ Acknowledgements.
Z. L. is partially supported by NSFC (22393913) and the robotic AI-Scientist platform of Chinese Academy of Sciences. R. C. acknowledges support from National Natural Science Foundation of China under Grant No. 12104328.
Appendix A More details of the multi-flavor GTEMPO method
The GTEMPO method is a fermionic analogy of the TEMPO method for bosonic impurity problems. In TEMPO, a normal MPS is used to represent the integrand of the discretized bosonic PI, e.g., the ADT, as it is a tensor of normal scalars. While in GTEMPO, we use GMPS to represent the ADT for fermionic PI, as the latter is a tensor of GVs (Grassmann tensor). The central ideal of GTEMPO to solve quantum impurity problems is to represent the ADT as a GMPS and then compute multi-time impurity correlations based on it. Implementation-wise, this is achieved in three major steps: (1) Discretization of the fermionic PI in Eq.(5), this is done by linearly discretizing the continuous Grassmann trajectories $\bm{a}_{p}(\tau)$, $\bar{\bm{a}}_{p}(\tau)$ into discrete GVs $\{a_{p,j}\}$ and $\{\bar{a}_{p,j}\}$ where $j$ labels the discrete time step, after which the ADT naturally becomes a Grassmann tensor; (2) Constructing the bare impurity dynamics part $\mathcal{K}[\bar{\bm{a}}\bm{a}]$ and each Feynman-Vernon IF $\mathcal{I}_{p}[\bar{\bm{a}}_{p}\bm{a}_{p}]$ as GMPSs; and (3) Multiplying $\mathcal{K}[\bar{\bm{a}}\bm{a}]$ and $\mathcal{I}_{p}[\bar{\bm{a}}_{p}\bm{a}_{p}]$ together to obtain the ADT using Eq.(11) and calculating multi-time impurity correlations based on it, e.g., Eq.(12). For step (1), there is a well-established scheme for the discretization, referred to as the the quasi-adiabatic propagator path integral (QuAPI) method [81, 82], for which one could also refer to Refs. [61, 55] for details. For step (2), essentially the only operation (besides the MPS bond truncation, which is a versatile operation that should be performed whenever the bond dimension of the MPS significantly grows) required is the multiplication between two GMPSs. For step (3), one needs to perform GMPS multiplication, as well as the integration of GMPS to obtain a scalar in the end. In addition, in step (3) we use the zipup algorithm to speed up the calculation of observables, which is a technique that evaluates the multiplication of several GMPSs on the fly instead of actually performing it, as a common practice in MPS algorithms to save memory and for computational efficiency. The only additional operation of GMPS required in multi-flavor GTEMPO, on top of vanilla GTEMPO, is the partial integration of a GMPS, which returns another GMPS with less sites instead of a scalar.
For make this work self-contained, we now discuss two key operations of GMPS, including GMPS multiplication, integration and partial integration of GMPS. We will mostly explain these operations based on their bosonic counterparts, i.e., the corresponding operations in TEMPO, as the latter are easier to understand, and then generalize to the fermionic case. One could also refer to Ref. [59, 73] for those details.
The first important operation for manipulating the GMPS is the element-wise product. In case of normal tensor, this operation is simply
$Z_{ijk\cdots}=X_{ijk\cdots}Y_{ijk\cdots},$ (20)
for two tensors $X$, $Y$ as an example. This operation can be straightforwardly extended to normal MPSs with two steps: (1) One could first perform a tensor product between the two MPSs. Assuming that the site tensors of these two GMPSs are written as $X_{\alpha_{i-1},\alpha_{i}}^{\sigma_{i}}$ and $Y_{\beta_{i-1},\beta_{i}}^{\sigma_{i}^{\prime}}$ where $\sigma_{i},\sigma_{i}^{\prime}$ are physical indices and $\alpha_{i},\beta_{i}$ are auxiliary indices, then the site tensor of the resulting MPS after the tensor product can be written as
$Z_{\alpha_{i-1}\beta_{i-1},\alpha_{i}\beta_{i}}^{\sigma_{i}\sigma_{i}^{\prime}}=X_{\alpha_{i-1},\alpha_{i}}^{\sigma_{i}}Y_{\beta_{i-1},\beta_{i}}^{\sigma_{i}^{\prime}};$ (21)
(2) Then one can perform the element-wise product for each pair of physical indices as
$W_{\alpha_{i-1}\beta_{i-1},\alpha_{i}\beta_{i}}^{\sigma_{i}}=Z_{\alpha_{i-1}\beta_{i-1},\alpha_{i}\beta_{i}}^{\sigma_{i}\sigma_{i}},$ (22)
with $W_{\alpha_{i-1}\beta_{i-1},\alpha_{i}\beta_{i}}^{\sigma_{i}}$ the site tensor of the final resulting MPS. We can see that the bond dimension of the resulting MPS is exactly the product of the bond dimensions of the two input MPSs, if MPS bond truncation is not performed. For two rank-$3$ Grassmann tensors denoted as $\mathcal{X}=\sum_{ijk}X_{ijk}\xi_{1}^{i}\xi_{2}^{j}\xi_{3}^{k}$ and $\mathcal{Y}=\sum_{ijk}Y_{ijk}\xi_{1}^{i}\xi_{2}^{j}\xi_{3}^{k}$, where $X$, $Y$ are coefficient tensors, and $\xi_{l}$ denotes the GV, the “element-wise” product is defined as
$\mathcal{Z}$ $=\mathcal{X}\mathcal{Y}=\sum_{ijk}X_{ijk}\xi_{1}^{i}\xi_{2}^{j}\xi_{3}^{k}\sum_{i^{\prime}j^{\prime}k^{\prime}}Y_{i^{\prime}j^{\prime}k^{\prime}}\xi_{1}^{i^{\prime}}\xi_{2}^{j^{\prime}}\xi_{3}^{k^{\prime}}$
$=\sum_{ijki^{\prime}j^{\prime}k^{\prime}}(-1)^{2i^{\prime}+k+j}(-1)^{j^{\prime}+k}X_{ijk}Y_{i^{\prime}j^{\prime}k^{\prime}}\xi_{1}^{i+i^{\prime}}\xi_{2}^{j+j^{\prime}}\xi_{3}^{k+k^{\prime}},$ (23)
where the two sign factors $(-1)^{2i^{\prime}+k+j}$ and $(-1)^{j^{\prime}+k}$ are due to the change of locations of the two GVs $\xi_{1}^{i^{\prime}}$ and $\xi_{1}^{j^{\prime}}$. The result of the element-wise product is still a rank-$3$ Grassmann tensor by recognizing that $\xi^{0}=1,\xi^{1}=\xi,\xi^{2}=0$ from the definition of GV. The GMPS multiplication inherits from the “element-wise” product between two Grassmann tensors, similar to the bosonic case, and the detailed implementation can be found in Ref. [59].
The second important operation for manipulating the GMPS is the integration and partial integration. To illustrate this operation we consider the integration of a pair of GVs, $\xi_{1}$ and $\xi_{2}$, from a rank-$3$ Grassmann tensor $\mathcal{X}=\sum_{ijk}X_{ijk}\xi_{1}^{i}\xi_{2}^{j}\xi_{3}^{k}$. This operation can be performed as
$\int\mathcal{D}[\xi_{1}\xi_{2}]\sum_{ijk}X_{ijk}\xi_{1}^{i}\xi_{2}^{j}\xi_{3}^{k}$
[…]
$=$ $\sum_{k}\left(\sum_{i}X_{iik}\right)\xi_{3}^{k},$ (24)
which results in a rank-$1$ Grassmann tensor in the end. In actual implementation of the Grassmann tensors, one only needs to stores the information of the coefficient tensor, and we can see that to perform a partial integration of the first pair of GVs one only needs to perform a partial trace of the first two indices of the coefficient tensor. In the general case that the pair of GVs been integrated out are not next to each other, one needs to first swap the GVs to put them in nearby positions, during which some sign factors may occur. Similarly, if we denote two nearby site tensors of a GMPS as $\sum_{\bar{\alpha}_{k+1},\sigma_{k+1},\alpha_{k}}A^{\sigma_{k+1}}_{\bar{\alpha}_{k+1},\alpha_{k}}\bar{\eta}_{k+1}^{\bar{\alpha}_{k+1}}\xi_{k+1}^{\sigma_{k+1}}\eta^{\alpha_{k}}$ and $\sum_{\bar{\alpha}_{k},\sigma_{k},\alpha_{k-1}}A^{\sigma_{k}}_{\bar{\alpha}_{k},\alpha_{k-1}}\bar{\eta}_{k}^{\bar{\alpha}_{k}}\xi_{k}^{\sigma_{k}}\eta^{\alpha_{k-1}}$, where $\bar{\eta}_{k},\eta_{k}$ denote the auxiliary GVs and $\xi_{k}$ denotes the physical GV, then we could integrate out these two physical GVs as follows:
[…]
$\left(\sum_{\sigma_{k+1},\bar{\alpha}_{k+1},\alpha_{k}}A^{\sigma_{k+1}}_{\bar{\alpha}_{k+1},\alpha_{k}}\bar{\eta}_{k+1}^{\bar{\alpha}_{k+1}}\xi_{k+1}^{\sigma_{k+1}}\eta_{k}^{\alpha_{k}}\right)\times$
$\left(\sum_{\sigma_{k},\bar{\alpha}_{k},\alpha_{k-1}}A^{\sigma_{k}}_{\bar{\alpha}_{k},\alpha_{k-1}}\bar{\eta}_{k}^{\bar{\alpha}_{k}}\xi_{k}^{\sigma_{k}}\eta_{k-1}^{\alpha_{k-1}}\right)$
$=$ $\sum_{\bar{\alpha}_{k+1},\alpha_{k-1}}\left(\sum_{\sigma_{k},\alpha_{k}}A^{\sigma_{k}}_{\bar{\alpha}_{k+1},\alpha_{k}}A^{\sigma_{k}}_{\alpha_{k},\alpha_{k-1}}\right)\bar{\eta}_{k+1}^{\bar{\alpha}_{k+1}}\eta_{k-1}^{\alpha_{k-1}},$ (25)
which boils down to the normal contraction of the two coefficient tensors. The result is a rank-$2$ Grassmann tensor without physical GVs, which can then be absorbed into nearby site tensors, and the net result is a new GMPS with the two sites corresponding to $\xi_{k+1}$ and $\xi_{k}$ missing compared to the original one. The partial integration, performed after the GMPS multiplication, is the major difference between our method to calculate the reduced ADT and the boundary MPS method to compress a 2D tensor network into 1D. As the information of the unused flavors are forgotten during this process, we observe that the bond dimension of resulting GMPS, i.e., $\chi_{2}$ can be greatly suppressed and grow much slower than the worst case $O(2^{n}\chi^{n})$, which is as expected. For full integration, one simply integrates out all the pairs of physical GVs, which would result in a single scalar.